### v1

@@ 4.1.1 Models performance across prompting strategies
This section analysis the translation performance of our first approach to DA-MSA translation. The reported scores are averaged across the three dialects represented in the test set. Detailed per-dialect results are available in Section  8. Table 7 summarises the mean CHrF++ scores for each model and prompting strategy.
@@
The two primary assumptions for ANOVA, normality and sphericity, were verified. We assessed the normality of the data distribution using the Shapiro-Wilk test, setting a significance level of 0.050.050.05 that must be exceeded to assume normality. To complement this analytical approach, we also visually inspected histograms and Quantile-Quantile (QQ) plots. Sphericity was evaluated using Mauchly’s Test of Sphericity, with the assumption being met if $p\leq 0.05$.
When the ANOVA indicated a statistically significant overall effect, we performed Bonferroni post-hoc tests (with 0.050.050.05 threshold for the $p$-value) to identify which specific pairs of prompting strategies differed significantly from one another. The results of these pairwise comparisons are presented in Table 6.
@@
Table 5: CHrF++ scores comparing models performance with zero-shot, CoT, Ara-TEaR and Few-shot prompting strategies for DA-MSA translation.
@@
Table 6: Statistically significance results where Strategy A achieves a statistically more significant score than Strategy B.
@@
Few-shot prompting emerges as the most effective strategy across the majority of tested models, including GPT-4o, LLaMA 3.3‑70B‑Instruct, LLaMA 3.1‑405B, Gemma‑27B, and DeepSeek‑V3. The consistent superiority of this method is likely due to the use of concrete translation examples tailored to each dialect, which enhances the models’ understanding of dialectal structures. By providing diverse and context-rich examples, few-shot prompting enables models with strong instruction-following capabilities to generalize more effectively and capture subtle linguistic nuances, hence achieving higher ChrF++ scores.
@@
Figure 1: Comparison of CHrF++ scores across models using Zero-Shot, Zero-Shot CoT, Few-Shot, and Ara-TEaR strategies for DA-MSA translation.
@@ 4.1.2 Dialect-identification impact on prompting performance
GPT-4o exhibited a moderate improvement in ChrF++ scores, increasing from 39.08 to 40.52. This gain reflects the model’s strong multilingual capabilities, likely stemming from extensive and diverse pre-training. In contrast, LLaMA 3.3-70B-Instruct demonstrated a more substantial improvement, with ChrF++ scores rising from 30.93 to 34.26. This suggests that LLaMA has limited capacity for implicit dialect identification and relies more heavily on explicit prompt structure to guide its translation decisions.
@@ 4.1.3 Cross-dialect performance variability
An analysis of per-dialect performance (presented in Section  8) reveals that Egyptian Arabic consistently yields the highest ChrF++ scores across nearly all models in the zero-shot setting, followed by Gulf Arabic and then Levantine. This outcome contrasts with established linguistic research, which identifies Levantine Arabic as morphologically and syntactically more similar to Modern Standard Arabic (MSA) Kwaik et al (2019). The observed disparity is likely attributable to the disproportionate representation of Egyptian Arabic in pre-training data, reflecting its widespread use in media and online content.
@@ 4.1.4 Computational Efficiency Analysis
Figure 2 presents the average time required to translate 200 Levantine sentences to MSA using Gemma2-9B via the LLAMA API, comparing the computational cost of each prompting strategy. Zero-shot (0.967s) and zero-shot CoT (0.973s) show minimal difference, indicating negligible overhead from reasoning prompts. Few-shot (0.989s) is slightly slower due to example processing.
@@
Figure 2: Comparison of average response times for zero-shot, chain-of-thought, few-shot, and Ara-TEaR strategies during DA-MSA translation.
4.2 Resource-efficient fine-tuning approach for DA–MSA translation
4.2.1 Experiment 0: Baseline Performance
@@
Although the model was not explicitly trained on dialect-specific data, it produced relatively consistent results across dialects. The narrow score range (3.8$\leq$ 3.8≤ 3.8 CHrF++) suggests minimal architectural bias toward any particular dialect, validating the model’s suitability as a baseline for our fine-tuning experiments.
4.2.2 Experiment 1: Hyperparameter Optimization
Experiment 1.1: Learning Rate Comparison
@@ 4.2.2 Experiment 1: Hyperparameter Optimization
Figure 3 illustrates the training loss curve for this experiment, using a smoothed line to emphasize the overall trend and overlook short-term fluctuations. The loss curves for both learning rates were nearly identical. Although the higher learning rate of 1e-4 produced a slightly better score, the learning rate of 5e-5 was selected due to its greater stability and lower risk of oscillations, which is especially important when scaling to larger datasets or longer training epochs in subsequent experiments.
@@
Experiment 1.2: Epoch Comparison
@@
The best configuration achieved a CHrF++ score of 43.44, representing a substantial improvement over the baseline score of 31.44 established in Experiment 0. Hyperparameter tuning resulted in a 38% improvement in CHrF++ over the baseline. Therefore, all subsequent experiments use a learning rate of 5e-5 and a single training epoch.
4.2.3 Experiment 2: Joint vs. Dialect-Specific Models
@@
The joint model outperformed the dialect-specific model by 10.6%, highlighting the benefits of cross-dialect training. As shown in Figure 4, there is substantial lexical overlap between Modern Standard Arabic (MSA) and the dialects: 18.7% for Levantine (LEV), 23.6% for Gulf (GLF), and 27.5% for Egyptian (EGY). This overlap, along with shared grammatical structures, enables better generalization in joint models. These results provide strong empirical support for employing joint modeling strategies in subsequent experiments.
Figure 4: Proportion of shared vocabulary between Modern Standard Arabic and Levantine, Gulf, and Egyptian dialects in the training dataset.
4.2.4 Experiment 3: Data Scaling Effects
@@
  • Model trained on D2: CHrF++ score of 49.88
  • Model trained on D1 (from Experiment 1): CHrF++ score of 43.44
Training on the expanded dataset (D2) yielded a 14.8% improvement in CHrF++ over the model trained on D1. A significant contributor to this gain was the inclusion of Dial2MSA’s social media data, which enhanced the model’s capacity to handle informal expressions and domain-specific linguistic patterns. These results underscore the value of both enlarging the training corpus and diversifying its domains.
4.2.5 Experiment 4: Data Augmentation
@@
Despite manual validation confirming the linguistic accuracy of the synthetic paraphrases, the augmented data introduced redundant lexical patterns. This redundancy led to overfitting in the Gemma-9B model, reducing its cross-dialect generalization.
@@
Due to computational constraints, this study could not investigate optimal synthetic-to-gold data ratios (e.g., 1:3), which may mitigate overfitting. Overall, the results emphasize that even high-quality synthetic data must be rigorously filtered, especially when training larger models, to avoid detrimental effects on performance.
4.2.6 Experiment 5: Model Architecture Comparison
@@
Table 7: Performance of fine-tuned and zero-shot models.
The results, as shown in Table 7, indicate that architectural efficiency and task-specific fine-tuning often outweigh raw parameter count. Gemma2-9B outperformed its larger counterpart, Gemma2-27B, by 1.39 CHrF++, pointing to diminishing returns from scale alone.
@@
In summary, the findings underscore the importance of model architecture and fine-tuning over sheer size. Gemma2-9B presents a compelling balance of accuracy and efficiency, making it well-suited for real-world deployment, particularly in resource-constrained environments.
Figure 5: Performance comparison between fine-tuned and zero-shot models on DA-MSA translation.
4.2.7 Experiment 6: Quantization Impact Analysis
@@
The 4-bit quantized model exhibited only a 0.9% decrease in CHrF++ score, a negligible reduction that confirms the viability of quantization for preserving translation quality. This supports the validity of using 4-bit quantization in earlier experiments, as the impact on performance remains within acceptable margins for practical use.
@@
4.2.8 Experiment 7: Active Learning Approach
@@
  • Baseline (fine-tuned on D2 dataset): CHrF++ score of 49.88
@@
Active learning led to a 1.5% reduction in CHrF++ score relative to the baseline, indicating that the sentence selection strategy was suboptimal. While longer sentences were assumed to offer richer syntactic structures, manual analysis revealed two key limitations:
@@
  • Limited Lexical Diversity: The selected sentences also exhibited low vocabulary variation, which restricted the model’s ability to generalize across dialects and reduced the effectiveness of the fine-tuning process.
@@ 5 Conclusions
This work focused on advancing DA-MSA machine translation through two approaches: training-free prompting and resource-efficient fine-tuning.
The training-free prompting experiments demonstrated that few-shot prompting consistently yielded strong performance across various LLMs, with GPT-4o achieving the highest CHrF++ scores. Our work also revealed that explicitly identifying the source dialect in prompts notably improved translation quality, particularly for models like LLAMA 3.3-70B-Instruct. Furthermore, our analysis showed that the prevalence of a dialect in pre-training data more strongly influenced translation performance than its linguistic similarity to MSA, with Egyptian Arabic consistently outperforming other dialects. The multi-stage Ara-TEaR method did not provide statistically significant gains and proved computationally intensive. The overall findings offer practical guidance and insights for prompt design in low-resource DA-MSA translation scenarios.
@@
  • Synthetic data augmentation introduced redundancy and negatively impacted performance, especially for larger models.
@@
In conclusion, our research provides a comprehensive framework for advancing DA-MSA machine translation, offering empirically supported strategies for both training-free and fine-tuning approaches. Our insights regarding prompt design, data strategies, model architecture, and efficiency considerations lay a strong foundation for future research and the development of robust and accessible DA-MSA translation systems in real-world low-resource applications.
@@ 6 Challenges and Limitations
Despite the progress achieved, our work faced several limitations. Computational constraints restricted the scope of our experiments, limiting hyperparameter tuning and preventing fine-tuning of larger models. Resource restrictions also limited data augmentation efforts, with synthetic paraphrasing leading to an unexpected 6% drop in CHrF++ scores, highlighting the risk of including synthetic data in training without careful filtering. The focus on three dialects (Levantine, Gulf, Egyptian) constrained generalizability, leaving the proposed translation techniques untested on varieties such as Maghrebi and Sudanese Arabic. Additionally, the overrepresentation of Egyptian Arabic in training data likely inflated performance for this dialect, prompting cautious interpretation of results.
Evaluation methodologies presented further challenges. Standard lexical metrics, such as SacreBLEU and CHrF++, inadequately captured meaning preservation, especially with dialect-specific expressions. Benchmarking our fine-tuned models against prior works proved difficult due to the absence of standardized test datasets and evaluation protocols, limiting the ability to make definitive performance claims.
These limitations highlight crucial areas for future work, including broader dialect coverage, improved evaluation metrics, and standardized benchmarking.
@@ 7 Further Work
Our future work could include investigating the performance of our models on a wider variety of dialects to identify any patterns and performance gaps. Refinements in data augmentation strategies are necessary to overcome the performance drop observed in our experiments. Future augmentation pipelines should integrate automatic filtering based on quality assessments to avoid training models on low-quality data. Given the current inability of language models to effectively back-translate from MSA to DA, alternative data augmentation techniques must be explored to enhance dataset diversity while maintaining high data quality.

### v2

@@ 4.1.1 Models performance across prompting strategies
This section analyses the translation performance of models in our first approach to DA-MSA translation. The reported scores are averaged across the three dialects represented in the test set. Detailed per-dialect results are available in Section 8. Table 7 summarizes the average chrF++ scores for each model and prompting strategy.
@@
The two primary assumptions for ANOVA, normality and sphericity, were verified. We assessed the normality of the data distribution using the Shapiro-Wilk test, setting a significance level of $0.05$ that must be exceeded to assume normality. To complement this analytical approach, we also visually inspected histograms and Quantile-Quantile (QQ) plots. Sphericity was evaluated using Mauchly’s Test of Sphericity, with the assumption being met if $p\leq 0.05$.
When the ANOVA indicated a statistically significant overall effect, we performed Bonferroni post-hoc tests (with $0.05$ threshold for the $p$-value) to identify which specific pairs of prompting strategies differed significantly from one another. The results of these pairwise comparisons are presented in Table 6.
@@
Table 5: Average chrF++ scores comparing models performance with zero-shot, CoT, Ara-TEaR and Few-shot prompting strategies for DA-MSA translation
@@
Table 6: Statistically significant results where Strategy A achieves a statistically more significant score than Strategy B
@@
Few-shot prompting emerges as the most effective strategy across the majority of tested models, including GPT-4o, LLaMA 3.3‑70B‑Instruct, LLaMA 3.1‑405B, Gemma‑27B, and DeepSeek-V3. The consistent superiority of this method is likely due to the use of concrete translation examples tailored to each dialect, which enhances the models’ understanding of dialectal structures. By providing diverse and context-rich examples, few-shot prompting enables models with strong instruction-following capabilities to generalize more effectively and capture subtle linguistic nuances, hence achieving higher chrF++ scores.
@@
Figure 1: Comparison of chrF++ scores across models using Zero-Shot, Zero-Shot CoT, Few-Shot, and Ara-TEaR strategies for DA-MSA translation
@@ 4.1.2 Dialect-identification impact on prompting performance
GPT-4o exhibited a moderate improvement in chrF++ scores, increasing from 39.08 to 40.52. This gain reflects the model’s strong multilingual capabilities, likely stemming from extensive and diverse pre-training. In contrast, LLaMA 3.3-70B-Instruct demonstrated a more substantial improvement, with chrF++ scores rising from 30.93 to 34.26. This suggests that LLaMA 3.3-70B-Instruct has limited capacity for implicit dialect identification and relies more heavily on explicit prompt structure to guide its translation decisions.
@@ 4.1.3 Cross-dialect performance variability
An analysis of per-dialect performance (presented in Section  8) reveals that Egyptian Arabic consistently yields the highest chrF++ scores across nearly all models in the zero-shot setting, followed by Gulf Arabic and then Levantine. This outcome contrasts with established linguistic research, which identifies Levantine Arabic as morphologically and syntactically more similar to MSA than many other dialects (Kwaik et al, 2018). The observed disparity is likely attributable to the disproportionate representation of Egyptian Arabic in pre-training data, reflecting its widespread use in media and online content.
@@ 4.1.4 Computational Efficiency Analysis
Figure 2 presents the average time required to translate 200 Levantine sentences to MSA using Gemma2-9B via the LLAMA API, comparing the model’s response time for each prompting strategy. Zero-shot (0.967s) and zero-shot CoT (0.973s) show minimal difference, indicating negligible overhead from reasoning prompts. Few-shot (0.989s) is slightly slower due to example processing.
@@
Figure 2: Comparison of average response times for zero-shot, chain-of-thought, few-shot, and Ara-TEaR strategies during DA-MSA translation
4.2 Evaluation of the Resource-efficient fine-tuning approach for DA–MSA translation
This section presents the results for the experiments outlined in Section  3.4.3.
4.2.1 Base: Baseline Performance (v0)
@@
Although the model was not explicitly trained on dialect-specific data, it produced relatively consistent results across dialects. The narrow score range ($\leq 3.8$ chrF++) suggests minimal architectural bias toward any particular dialect, validating the model’s suitability as a baseline for our fine-tuning experiments.
4.2.2 HypperParam: Hyperparameter Optimization (v1)
HypperParam-LR: Learning Rate Comparison
@@ 4.2.2 HypperParam: Hyperparameter Optimization (v1)
Figure 3 illustrates the training loss curve for this experiment, using a smoothed line to emphasize the overall trend and overlook short-term fluctuations. The loss curves for both learning rates were nearly identical. Although the higher learning rate of 1e-4 produced a slightly higher score, the learning rate of 5e-5 was selected due to its likely greater stability and lower risk of oscillations, which is especially important when scaling to larger datasets or longer training epochs in subsequent experiments.
@@
HypperParam-Epoch: Epoch Comparison
@@
The best configuration achieved a chrF++ score of 43.44, representing a substantial improvement over the baseline score of 31.44 established in the Base Experiment. Model fine-tuning and hyperparameter-tuning resulted in a 38% improvement in chrF++ over the baseline. Therefore, all subsequent experiments use a learning rate of 5e-5 and a single training epoch.
4.2.3 Join-vs-Spec: Joint vs. Dialect-Specific Models (v2)
@@
The joint model outperformed the dialect-specific model by 10.6%, highlighting the benefits of cross-dialect training. As shown in Figure 4, there is substantial lexical overlap between MSA and the dialects: 18.7% for Levantine (LEV), 23.6% for Gulf (GLF), and 27.5% for Egyptian (EGY). This overlap, along with shared grammatical structures, enables better generalization in joint models. These results provide strong empirical support for employing a joint modeling strategy in subsequent experiments.
Figure 4: Proportion of shared vocabulary between Modern Standard Arabic and Levantine, Gulf, and Egyptian dialects in the training dataset
4.2.4 DataScale: Data Scaling Effect (v3)
@@
  • Model trained on Extended-Gold dataset: chrF++ score of 49.88
  • Model trained on the Minimal-Gold dataset (from HypperParam experiment): chrF++ score of 43.44
Training on the Extended-Gold dataset yielded a 14.8% improvement in chrF++ over the model trained on the Minimal-Gold dataset. A significant contributor to this gain was the inclusion of Dial2MSA-Verified dataset (which has a social media domain). Hence, enhancing the model’s capacity to handle informal expressions and domain-specific linguistic patterns. These results underscore the value of both enlarging the training corpus and diversifying its domain.
4.2.5 Augmentation: Data Augmentation Effect (v4)
@@
Despite manual validation confirming the linguistic accuracy of the synthetic paraphrases, the augmented data has likely introduced redundant lexical patterns. This redundancy led to overfitting in the Gemma-9B model, reducing its cross-dialect generalization.
@@
Due to computational constraints, this study could not investigate optimal synthetic-to-gold data ratios (e.g., 1:3), which may mitigate overfitting. Overall, the results emphasize that even high-quality synthetic data must be rigorously filtered, especially when training larger models, to avoid detrimental effects on performance. While our synthetic dataset did not improve the performance, we made it publicly accessible to allow for further investigation and potential use in other Arabic NLP tasks.
4.2.6 Architecture: Model Architecture Comparison (v5)
@@
Table 7: Performance of fine-tuned and zero-shot models
The results, as shown in Table 7, indicate that task-specific fine-tuning often outweigh raw parameter count. Gemma2-9B outperformed its larger counterpart, Gemma2-27B, by 1.39 chrF++, pointing to potential diminishing returns from scale alone without improving the training strategy and training dataset.
@@
In summary, the findings underscore the importance of the model architecture choice and task-specific fine-tuning over sheer size. Gemma2-9B presents a compelling balance of accuracy and efficiency, making it well-suited for real-world deployment, particularly in resource-constrained environments. Figure 5 provides an alternative visual presentation of the results.
Figure 5: Performance comparison between fine-tuned and zero-shot models on the Large-Test dataset
4.2.7 Quantiz: Quantization Impact (v6)
@@
The 4-bit quantized model exhibited only a 0.9% decrease in chrF++ score, a negligible reduction that confirms the viability of quantization while preserving translation quality. This supports the validity of using 4-bit quantization in earlier experiments, as the impact on performance remains within acceptable margins for practical use.
@@
4.2.8 ActiveLearn: Active Learning Approach (v7)
@@
  • Baseline (fine-tuned on Extended-Gold dataset): chrF++ score of 49.88
@@
Active learning led to a 1.5% reduction in chrF++ score relative to the baseline, indicating that the sentence selection strategy was suboptimal. While longer sentences were assumed to offer richer syntactic structures, we suspect two key limitations after performing manual analysis of the sentences selected for active learning:
@@
  • Limited Lexical Diversity: The selected sentences exhibited low vocabulary variation, which restricted the model’s ability to generalize across dialects and reduced the effectiveness of the fine-tuning process.
@@
4.3 Comparative evaluation between model fine-tuning and training-free prompting for DA–MSA translation
While the evaluation of training-free prompting techniques in Section 4.1 was conducted on the Small-Test dataset due to the prohibitive cost of model prompting, the larger (Large-Test) dataset was used for evaluating fine-tuned models in Section 4.2. To enable direct comparison between the two approaches, in this section we evaluate all developed solutions (both fine-tuned and prompted models) on the Small-Test dataset. Figure 6 presents a summary of the results.
The results demonstrate that fine-tuning generally outperforms zero-shot prompting, with Gemma2-9B achieving the highest chrF++ score (49.17), surpassing even GPT-4o (42.18) despite its vastly smaller size. Gemma2-27B also performs strongly (44.56) but trails behind the 9B variant, suggesting diminishing returns from scaling under limited or noisy data. Fine-tuned LLaMA models achieve moderate gains, while prompted models cluster in the mid-30s to low-40s range. Overall, the findings highlight that targeted fine-tuning of mid-sized models provides the best balance of accuracy and efficiency for DA–MSA translation.
Figure 6: Comparison of chrF++ scores across different fine-tuned and prompted models on the Small-Test dataset
@@ 5 Conclusions
This work focused on evaluating and advancing DA-MSA machine translation through two approaches: training-free prompting and resource-efficient fine-tuning.
The training-free prompting experiments demonstrated that few-shot prompting consistently yielded strong performance across various LLMs, with GPT-4o achieving the highest chrF++ scores. Our work also revealed that explicitly identifying the source dialect in prompts notably improved translation quality, particularly for models like LLAMA 3.3-70B-Instruct. Furthermore, our analysis showed that the prevalence of a dialect in pre-training data more strongly influenced translation performance than its linguistic similarity to MSA, with Egyptian Arabic consistently outperforming other dialects. Our multi-stage Ara-TEaR method did not provide statistically significant gains and proved computationally expensive. The overall findings offer practical guidance and insights for prompt design in low-resource DA-MSA translation scenarios.
@@
  • Synthetic data augmentation without careful filtering negatively impacted performance, especially for larger models.
@@ 6 Challenges and Limitations
Despite the progress achieved, our work faced several limitations. Computational constraints restricted the scope of our experiments, limiting hyperparameter tuning and preventing fine-tuning of larger models. Resource restrictions also limited our data augmentation efforts. The focus on three dialects (Levantine, Gulf, Egyptian) constrained generalization, leaving the proposed translation techniques untested on varieties such as Maghrebi and Sudanese Arabic. Additionally, the overrepresentation of Egyptian Arabic in training data likely inflated performance for this dialect, prompting cautious interpretation of results.
Evaluation methodologies presented further challenges. Standard lexical metrics, such as SacreBLEU and CHrF++, are likely incapable of fully capturing meaning preservation in translations , especially with dialect-specific expressions. Comparing our fine-tuned models against prior works proved difficult due to the absence of public and standardized benchmark datasets and evaluation protocols, limiting the ability to make definitive performance claims.
These limitations highlight crucial areas for future work, including broader dialect coverage, improved evaluation metrics, and standardized benchmark development.
@@ 7 Further Work
Our future work could include investigating the performance of our models on a wider variety of dialects to identify any patterns and performance gaps. Refinements in data augmentation strategies are necessary to overcome the performance drop observed in our experiment. Future augmentation pipelines should integrate automatic filtering based on quality assessments to avoid training models on low-quality data. Given the current inability of language models to effectively back-translate from MSA to DA, alternative data augmentation techniques must be explored to enhance dataset diversity while maintaining high data quality.